Operations Report – January 10th

Operations Report

Crew 219 Operations Report 10-JAN-2020

Sol: 5

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 14

From what time last night: 17:00

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 84%

10:30 =100%, 11:00 =99%, 11:30 = 98%, 12:00 =97%, 12:30= 96%, 13:00 =
95%, 13:30 = 93%, 14:00 = 91%, 14:30 = 90%, 15:00 = 89%, 15:30 = 88%,
16:00 = 87%, 16:30 = 87%, 17:00 = 84%

Diesel Reading – 70%

Station Propane Reading- 45%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147538,3

Water (static tank) (gal) – 468

Static to loft Pump used – Yes

Water in Green Hab (gal): 188.65

Water in Science Dome (gal): 0

Toilet Tank Emptied – No

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- No

Starting Hours:

Beginning Charge:

Ending Hours:

Ending Charge:

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: No

Starting Hours:

Beginning Charge:

Ending Hours:

Ending Charge:

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Used by Outpost Team to go to the storage
room, get water and
supplies for Outpost and Hab.

CrewCar used and why, where? Off-Campus, still on the workshop for repair.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Defective radios have been placed inside
the RAM for Outpost evaluation. The repairs for suit #3 was tested on
EVA 10 today and performed perfectly. All two-piece helmets are
missing one O-Ring for their connectors.

Summary of Hab: Pipe froze again today. The heating coil is still
functioning and the pump is still operating properly. The freezing
seems to be in between the exterior wall and the interior wall leading
to the washroom. This was solved using a combined method of warm water
through the feedline, a hair dryer heating the interior. The
thermostat in the was room read lowest temperature at -2.0°C during
the night. Outpost did a repair today to the exterior section of the
pipe. On the 9th, solar panels were inspected for any damage. All the
system and structure are nominal and fully operational. Heat tape for
the pipelines from the static tank to the Hab is being rearranged in
order to completely fix the frozen pipes issue. Furnace filter was
also changed today. We received Teflon tape, toilet paper, and pipe
insulation.

Summary of Science Dome operations: Temperatures are hovering around 7
°C during the day.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: 11 O-Rings (7
replacements and 4 Bench stock) and new transformer for radio charging
unit (input=120VAC 60Hz 9W, Output=9V DC 300 mA)